# Modeling Advection-Dominated Flows with Space-Local Reduced-Order Models

### v2

@@ 3.2 Space-local POD
In space-local proper orthogonal decomposition (L-POD) we take a different approach from G-POD. We start off by assuming a uniform finite difference discretization for the discrete FOM solution $u_{h}(x,t)$ [^1] If $u_{h}(x,t)$ would stem from a simulation on an unstructured grid, one could potentially first project the discrete solution on a uniform grid, after which the presented methodology could still be applied. We consider this outside the scope of this research.. Once the solution is represented on this grid we subdivide the domain $\Omega$ into $I$ non-overlapping subdomains $\Omega_{i}=[\alpha_{i},\beta_{i})$, $i=1\ldots I$, such that $\alpha_{i}<\beta_{i}=\alpha_{i+1}<\beta_{i+1}$. Later in this section these subdomains will be used to construct the L-POD basis. Furthermore, we assume each of these subdomains contains exactly $J$ grid points. This means the total number of grid points $N$ has to equal the product $N=I\cdot J$. This is essential for our methodology to work, as for the L-POD procedure each subdomain is treated equally in the snapshot matrix which is only possible if they contain the same number of grid points. This helps us obtain a basis that generalizes better outside the training data for advection-dominated problems, where the dynamics are similar throughout the domain. Such as for a traveling wave. This is fundamentally different from the approach presented in [25], where each subdomain has its own POD basis. The latter is more suitable for problems where the dynamics differ throughout the domain, such that the local bases can be specialized to these dynamics. If $N/I$ is not an integer one could possibly resolve this by projecting the FOM solution on a compatible grid of $M$ grid points for which $M/I$ is an integer.
@@
After obtaining $\mathbf{U}(t)$ at different points in time the snapshot matrix is constructed as follows
@@
In this way each subdomain is treated equally in the snapshot matrix. This is what differentiates our work from [25], namely the snapshot matrix is built combining all of the subdomains. Using a common snapshot matrix allows us to obtain a space-local POD basis that generalizes well outside the training data. The way we divide the domain in subdomains is decided by what results in an L-POD basis that generalizes best on a validation data set, see section 5.2.
Note that this matrix has fewer rows, but more columns than the G-POD snapshot matrix, see (11). This makes the SVD cheaper to compute for a large number of snapshots $s$ [29]. As in the global case, we use a SVD of the snapshot matrix to obtain a truncated basis $\hat{\boldsymbol{\Gamma}}\in\mathbb{R}^{J\times q}$ with $q<J$ from the left-singular vector. In terms of the local basis $\{\chi_{ij}\}$ the POD basis is written as
@@
analogous to (18). In this basis the approximated solution $u_{\ell}$ is obtained as
@@
This matrix, containing non-overlapping blocks, is constructed such that multiplying by its transpose projects $\mathbf{u}$ onto the L-POD basis. For the mapping from $\mathbf{a}_{\ell}\in\mathbb{R}^{Iq}$ to $\mathbf{A}\in\mathbb{R}^{I\times q}$ we follow the same convention as for $\mathbf{u}$ and $\mathbf{U}$. The effective number of basis functions for L-POD is $Iq=r$ with $q<J$. This basis is orthonormal, i.e. $\boldsymbol{\Gamma}^{T}\boldsymbol{\Gamma}=\mathbf{I}$. The sparsity of $\boldsymbol{\Gamma}$, as opposed to $\boldsymbol{\Phi}$, is what yields a ROM which is cheaper to evaluate than a G-POD-based ROM [25].
An example of an L-POD approximation $u_{\ell}$ to a Gaussian wave is shown in Figure 2. For the L-POD we used $I=10$ subdomains with $q=6$ basis functions per subdomain. The training data used to obtain the basis is discussed in the results section (section 5), see Figure 4. Although the L-POD approximation is accurate in most of the domain, some oscillations appear near the boundaries. In addition, the approximation is not guaranteed to be smooth on the edges of the subdomains, see Figure 2. Discontinuities appear when transitioning from one subdomain to the next. In section 5 we will show that these discontinuities tend to grow increasingly severe when using the L-POD basis to construct a ROM, see Figure 7. To remedy this issue, we introduce space-local POD with overlapping subdomains in the next section.
@@ 3.3 Local POD with overlapping subdomains
In finite element discretizations similar discontinuities are resolved by imposing continuity of the approximation, while using a local basis. In this work we aim to achieve the same by introducing a novel space-local POD formulation with overlapping subdomains. We will refer this approach as LO-POD. To start off, we subdivide the domain $\Omega$ into $I$ overlapping subdomains $\Omega_{i}=[\alpha_{i},\beta_{i})$. This means $\beta_{i-2}=\alpha_{i}<\beta_{i-1}=\alpha_{i+1}<\beta_{i}=\alpha_{i+2}$. Once again, each subdomain contains $J$ grid points. Note that due to the overlap each grid point is now located in two subdomains. This means $\frac{I\cdot J}{2}=N$, as opposed to $I\cdot J=N$ for L-POD.
In order to obtain a smooth approximation $u_{\ell}$ we require the LO-POD basis to smoothly decay to zero on the edge of the subdomain. This is enforced through a post-processing step of the local snapshot matrix given by (21). For this purpose we introduce a kernel $k_{i}(x)$ to divide the solution between the subdomains. This places the following constraint on the kernels:
@@
such that this set of functions forms a partition of unity. Here we propose the following kernel
@@
which is chosen as it smoothly decays to zero at the subdomain boundaries. This approach is inspired by the partition of unity method [30]. Here the right half of each subdomain is overlapped by the subdomain on the right and the left half by the subdomain on the left. A visualization of the kernels is displayed in Figure 3.
@@
These integrals can simply be approximated using the midpoint rule for integration [31]. The remaining procedure stays the same as for L-POD, i.e. we build the snapshot matrix in (21) and carry out a SVD to obtain $\hat{\boldsymbol{\Gamma}}$. However, what changed is that the blocks in $\boldsymbol{\Gamma}$, see (25), are now overlapping. This means the basis is no longer orthonormal, i.e. $\boldsymbol{\Gamma}^{T}\boldsymbol{\Gamma}\neq\mathbf{I}$. Due to the non-orthonormality, the coefficients $\mathbf{a}_{\ell}$ for the expansion in (23) are obtained by solving the following linear system
@@
as opposed to (24). Obtaining the coefficients from the FOM solution is therefore more expensive. In addition, evaluating the resulting ROM requires solving a linear system, as will be discussed in section 4. This makes the LO-POD ROM more expensive to evaluate than its non-overlapping counterpart L-POD, (see sections 4.2 and 5.5). However, looking at Figure 2, we find that using the LO-POD results in a much smoother approximation of the Gaussian wave. Note that the parameters are kept the same, i.e. $I=10$ and $q=6$. In section 5.2 the associated error will be quantified more precisely.
@@ 4.1 Projection on reduced basis
Consider again $\mathbf{c}$, the state vector of the full-order model, and $\boldsymbol{\Phi}\in\mathbb{R}^{N\times r}$ a reduced basis where $r<N$, obtained from either G-POD, L-POD, or LO-POD. For the space-local approaches this operator was referred to as $\boldsymbol{\Gamma}$, see section 3.2. The corresponding ‘Gram’ matrix $\mathbf{S}\in\mathbb{R}^{r\times r}$ of this basis is computed as
@@
Note that for both G-POD and L-POD computing the inverse of $\mathbf{S}$ is trivial, as it is simply the identity. This comes from the fact that the basis is orthonormal. However, for LO-POD the basis is non-orthogonal, which makes computing its inverse less trivial. As stated earlier, this is a downside to LO-POD and increases its computational cost. It is quite straightforward to see that $\mathbf{P}$ is idempotent as $\mathbf{P}^{2}=\mathbf{P}$. The POD coefficient vector is obtained as
@@ 4.2 Galerkin projection
Having obtained a reduced basis, we construct a ROM for the advection equation as follows [7, 8, 5, 6]: based on (6) we define the residual $\mathbf{r}\in\mathbb{R}^{N}$ as
@@
where we replaced $\mathbf{u}$ by the POD approximation $\mathbf{u}_{r}=\boldsymbol{\Phi}\mathbf{a}$. As $\mathbf{u}$ comes from the finite difference discretization of the advection equation we have $\mathbf{c}(t)=\mathbf{u}(t)$. Next, we carry out the Galerkin projection of the residual on the POD basis, according to (32), and set this to zero:
@@
Note that we introduced $\mathbf{a}^{\text{ROM}}$ here as the POD state vector predicted by ROM. This is typically not equal to the true $\mathbf{a}$, see (32), past $t=0$. Going from the FOM to the ROM we reduced the DOF in the system from $N$ to $r$. In the G-POD case $\mathbf{A}$ is typically dense, whereas in the L-POD/LO-POD it is sparse. The sparsity decreases the cost of evaluating the ROM [25]. The amount of nonzero entries in the G-POD ROM operator scales with $\mathcal{O}(r^{2})$. For L-POD it scales with $\mathcal{O}(((n+1)q)^{2}I)$, where $n$ is the number of neighbors per subdomain. This accounts for the interaction between the subdomains. For LO-POD the neighbors of the neighbors have to be included in the interactions due to the overlap. This results in a scaling of $\mathcal{O}(((\tilde{n}+1)q)^{2}I)$, where $\tilde{n}$ is the number of subdomains that can be reached by crossing at most one subdomain starting from one of the subdomains. For 1D problems $n=2$ and $\tilde{n}=4$. For a large number of subdomains $I$ and a small number of POD basis functions per subdomain $q$, with the total number of POD modes being $I\cdot q=r$, this scales more favorably than G-POD. This allows us to include a larger number of basis functions in the POD basis without sacrificing computational efficiency. Hyperreduction techniques, including energy-conserving ones, can also be employed to further sparsen the ROM operators [32, 33, 17].
@@ 4.3 Energy conservation of the ROM
In order to ensure stability of the ROMs, we aim to mimic the energy conservation property of the FOM, see (8). To investigate if the constructed ROMs satisfy this property we define the ROM energy $E_{r}^{\text{ROM}}$ as
@@
employing the product rule of differentiation. From the final expression we conclude that a ROM based on Galerkin projection inherits energy conservation and stability from the FOM. This is true for both orthogonal and non-orthogonal projections. This means that not only the G-POD ROM inherits this property, as shown in [9], but also the newly introduced space-local approaches L-POD and LO-POD.
@@ 5.1 Test case setup
To construct a ROM we first require data from the FOM. As stated earlier, we use the finite difference discretization of the advection equation, detailed in section 2.2, for this purpose. The system is simulated on a periodic domain $\Omega=[0,2\pi)$ discretized with $N=1000$ grid points for $t=[0,5]$ and constant $c=1$. A time step size of $\Delta t=0.01$ is used for the time integration, using a RK4 scheme. This is the largest time step size which still yielded stable simulations. Snapshot data for the ROM construction is collected in the interval $t=[0,1]$. We refer to this as the training data. Data generated in the interval $t=(1,2]$ will be referred to as validation data. This data is used to evaluate the generalization of the POD basis. The remaining part of the simulation data, $t=(2,5]$, will be used to evaluate the extrapolation capabilities of the ROMs. As an initial condition we use a Gaussian wave centered around $x=\frac{1}{4}\pi$, namely
@@
Figure 4: Reference simulation of a Gaussian wave being advected throughout the domain. White bars separate training data from validation data and validation data from extrapolation data, respectively.
For the ROMs we consider the three bases discussed in this work: the global POD basis G-POD, the space-local POD basis L-POD, and the space-local POD basis with overlapping subdomains LO-POD. For the space-local approaches the domain is subdivided into $I$ subdomains. For the local basis $\{\chi_{ij}\}$ we use $J$ box functions contained within each subdomain, such that $IJ=N$. In this way the local basis aligns with the FOM finite difference basis, see (10). The integrals in the LO-POD post-processing step, equation (28), are approximated using the midpoint rule for integration [31]. For the ROM time integration we use the same RK4 scheme as for the FOM.
@@
$\underbrace{\frac{\mathbf{u}^{\text{ROM}}_{r}(t)-\mathbf{u}^{\text{FOM}}(t)}{||\mathbf{u}^{\text{FOM}}(t)||_{2}}}_{:=\text{solution error}}=\underbrace{\frac{\mathbf{u}_{r}^{\text{ROM}}(t)-\mathbf{P}\mathbf{u}^{\text{FOM}}(t)}{||\mathbf{u}^{\text{FOM}}(t)||_{2}}}_{:=\text{ROM error}}+\underbrace{\frac{\mathbf{P}\mathbf{u}^{\text{FOM}}(t)-\mathbf{u}^{\text{FOM}}(t)}{||\mathbf{u}^{\text{FOM}}(t)||_{2}}}_{:=\text{projection error}},$ (41)
where $\mathbf{P}$ projects the solution on the POD basis, see (31). This difference will be referred to as the solution error. In (41) this error is decomposed as the sum of the ROM error (the error made by the ROM during the time integration) and the projection error (the error made by the reduced basis approximation). Our implementation, in Julia [34], of the introduced methodologies and experiments are freely available on Github, see https://github.com/tobyvg/local_POD_overlap.jl.

### v3

@@ 3.2 Space-local POD
In space-local proper orthogonal decomposition (L-POD) we take a different approach from G-POD. We start by assuming a uniform finite difference discretization for the discrete FOM solution $u_{h}(x,t)$ [^1] If $u_{h}(x,t)$ would stem from a simulation on an unstructured grid, one could potentially first project the discrete solution on a uniform grid, after which the presented methodology could still be applied. We consider this outside the scope of this research.. Once the solution is represented on this grid, we subdivide the domain $\Omega$ into $I$ non-overlapping subdomains $\Omega_{i}=[\alpha_{i},\beta_{i})$, $i=1\ldots I$, such that $\alpha_{i}<\beta_{i}=\alpha_{i+1}<\beta_{i+1}$. Later in this section, these subdomains will be used to construct the L-POD basis. Furthermore, we assume each of these subdomains contains exactly $J$ grid points. This means the total number of grid points $N$ has to equal the product $N=I\cdot J$. This is essential for our methodology to work, as for the L-POD procedure, each subdomain is treated equally in the snapshot matrix, which is only possible if they contain the same number of grid points. This helps us obtain a basis that generalizes better outside the training data for advection-dominated problems, where the dynamics are similar throughout the domain. This approach is fundamentally different from what is presented in [farhat_local_POD], where each subdomain has its own POD basis. The latter is more suitable for problems where the dynamics differ throughout the domain, such that the local bases can be specialized to these dynamics. If $N/I$ is not an integer, one could possibly resolve this by projecting the FOM solution on a compatible grid of $M$ grid points for which $M/I$ is an integer.
@@
After obtaining $\mathbf{U}(t)$ at different points in time, the snapshot matrix is constructed as follows
@@
In this way, each subdomain is treated equally in the snapshot matrix. This is what differentiates our work from [farhat_local_POD] and similar to what is done [CHUNG2024DDROM]. A schematic representation of this is shown in Figure 2, where the G-POD snapshot matrix $\mathbf{X}$ is reshaped into the L-POD snapshot matrix $\mathbf{X}_{\ell}$.
[…]
Figure 2: Schematic representation of the G-POD snapshot matrix $\mathbf{X}$ being reshaped into the L-POD snapshot matrix $\mathbf{X}_{\ell}$ for $I=3$.
Using a single common snapshot matrix allows us to obtain a space-local POD basis that generalizes well outside the training data. The way we divide the domain into subdomains is decided by what results in an L-POD basis that generalizes best on a validation data set, see section 5.2.
Note that this matrix has fewer rows but more columns than the G-POD snapshot matrix, see (11). This makes the SVD cheaper to compute for a large number of snapshots $s$ [SVD_complexity]. As in the global case, we use a SVD of the snapshot matrix to obtain a truncated basis $\hat{\boldsymbol{\Gamma}}\in\mathbb{R}^{J\times q}$ with $q<J$ from the left-singular vector. In terms of the local basis $\{\chi_{ij}\}$ the POD basis is written as
@@
analogous to (18). In this basis, the approximated solution $u_{\ell}$ is obtained as
@@
This matrix, containing non-overlapping blocks, is constructed such that multiplying by its transpose projects $\mathbf{u}$ onto the L-POD basis. For the mapping from $\mathbf{a}_{\ell}\in\mathbb{R}^{Iq}$ to $\mathbf{A}\in\mathbb{R}^{I\times q}$ we follow the same convention as for $\mathbf{u}$ and $\mathbf{U}$. The effective number of basis functions for L-POD is $Iq=r$ with $q<J$. This basis is orthonormal, i.e. $\boldsymbol{\Gamma}^{T}\boldsymbol{\Gamma}=\mathbf{I}$. The sparsity of $\boldsymbol{\Gamma}$, as opposed to $\boldsymbol{\Phi}$, is what yields a ROM which is cheaper to evaluate than a G-POD-based ROM [farhat_local_POD].
An example of an L-POD approximation $u_{\ell}$ to a Gaussian wave is shown in Figure 3. For the L-POD we used $I=10$ subdomains with $q=6$ basis functions per subdomain. The training data used to obtain the basis are discussed in the results section (Section 5), as shown in Figure 5. Although the L-POD approximation is accurate in most of the domain, some oscillations appear near the boundaries. In addition, the approximation is not guaranteed to be smooth on the edges of the subdomains, see Figure 3. Discontinuities appear when transitioning from one subdomain to the next. In Section 5, we will show that these discontinuities tend to grow increasingly severe when using the L-POD basis to construct a ROM, see Figure 8. To remedy this issue, we introduce space-local POD with overlapping subdomains in the next section.
@@ 3.3 Local POD with overlapping subdomains
In finite element discretizations, similar discontinuities are resolved by imposing continuity of the approximation, while using a local basis. In this work, we aim to achieve the same by introducing a novel space-local POD formulation with overlapping subdomains. We will refer to this approach as LO-POD. To start off, we subdivide the domain $\Omega$ into $I$ overlapping subdomains $\Omega_{i}=[\alpha_{i},\beta_{i})$. This means $\beta_{i-2}=\alpha_{i}<\beta_{i-1}=\alpha_{i+1}<\beta_{i}=\alpha_{i+2}$. Once again, each subdomain contains $J$ grid points. Note that due to the overlap, each grid point is now located in two subdomains. This means $\frac{I\cdot J}{2}=N$, as opposed to $I\cdot J=N$ for L-POD.
To obtain a smooth approximation $u_{\ell}$, we require the LO-POD basis to smoothly decay to zero on the edge of the subdomain. This is enforced through a post-processing step of the local snapshot matrix given by (21). For this purpose, we introduce a kernel $k_{i}(x)$ to divide the solution between the subdomains. This places the following constraint on the kernels:
@@
such that this set of functions forms a partition of unity. Here, we propose the following kernel
@@
which is chosen as it smoothly decays to zero at the subdomain boundaries. This approach is inspired by the partition of unity method [PUM]. Here, the right half of each subdomain is overlapped by the subdomain to its right, and the left half by the subdomain to its left. A visualization of the kernels is displayed in Figure 4.
@@
These integrals can simply be approximated using the midpoint rule for integration [midpoint_dragomir1998some]. The remaining procedure stays the same as for L-POD, i.e., we build the snapshot matrix in (21) and carry out a SVD to obtain $\hat{\boldsymbol{\Gamma}}$. However, what changed is that the blocks in $\boldsymbol{\Gamma}$, see (25), are now overlapping. This means the basis is no longer orthonormal, i.e. $\boldsymbol{\Gamma}^{T}\boldsymbol{\Gamma}\neq\mathbf{I}$. Due to the non-orthonormality, the coefficients $\mathbf{a}_{\ell}$ for the expansion in (23) are obtained by solving the following linear system
@@
as opposed to (24). Obtaining the coefficients from the FOM solution is therefore more expensive. In addition, evaluating the resulting ROM requires solving a linear system, as will be discussed in Section 4. This makes the LO-POD ROM more expensive to evaluate than its non-overlapping counterpart L-POD, see Sections 4.2 and 5.5. However, looking at Figure 3, we find that using the LO-POD results in a much smoother approximation of the Gaussian wave. Note that the parameters are kept the same, i.e., $I=10$ and $q=6$. In Section 5.2, the associated error will be quantified more precisely.
@@ 4.1 Projection on reduced basis
Consider again $\mathbf{c}$, the state vector of the full-order model, and $\boldsymbol{\Phi}\in\mathbb{R}^{N\times r}$ a reduced basis where $r<N$, obtained from either G-POD, L-POD, or LO-POD. For the space-local approaches, this operator was referred to as $\boldsymbol{\Gamma}$, see Section 3.2. The corresponding ‘Gram’ matrix $\mathbf{S}\in\mathbb{R}^{r\times r}$ of this basis is computed as
@@
Note that for both G-POD and L-POD, computing the inverse of $\mathbf{S}$ is trivial, as it is simply the identity. This is because the basis is orthonormal. However, for LO-POD the basis is non-orthogonal, which makes computing its inverse less trivial. As stated earlier, this is a downside to LO-POD and increases its computational cost. It is quite straightforward to see that $\mathbf{P}$ is idempotent as $\mathbf{P}^{2}=\mathbf{P}$. The POD coefficient vector is obtained as
@@ 4.2 Galerkin projection
Having obtained a reduced basis, we construct a ROM for the advection equation as follows [POD1, POD2, brunton2019data, projection_ROMs_survey]: based on (6) we define the residual $\mathbf{r}\in\mathbb{R}^{N}$ as
@@
where we replaced $\mathbf{u}$ by the POD approximation $\mathbf{u}_{r}=\boldsymbol{\Phi}\mathbf{a}$. As $\mathbf{u}$ comes from the finite difference discretization of the advection equation, we have $\mathbf{c}(t)=\mathbf{u}(t)$. Next, we carry out the Galerkin projection of the residual on the POD basis, according to (32), and set this to zero:
@@
Note that we introduced $\mathbf{a}^{\text{ROM}}$ here as the POD state vector predicted by ROM. This is typically not equal to the true $\mathbf{a}$, see (32), past $t=0$. Going from the FOM to the ROM, we reduced the DOF in the system from $N$ to $r$. In the G-POD case $\mathbf{A}$ is typically dense, whereas in the L-POD/LO-POD it is sparse. The sparsity decreases the cost of evaluating the ROM [farhat_local_POD]. The amount of nonzero entries in the G-POD ROM operator scales with $\mathcal{O}(r^{2})$. For L-POD it scales with $\mathcal{O}(((n+1)q)^{2}I)$, where $n$ is the number of neighbors per subdomain. This accounts for the interaction between the subdomains. For LO-POD, the neighbors of the neighbors have to be included in the interactions due to the overlap. This results in a scaling of $\mathcal{O}(((\tilde{n}+1)q)^{2}I)$, where $\tilde{n}$ is the number of subdomains that can be reached by crossing at most one subdomain starting from one of the subdomains. For 1D problems $n=2$ and $\tilde{n}=4$. For a large number of subdomains $I$ and a small number of POD basis functions per subdomain $q$, with the total number of POD modes being $I\cdot q=r$, this scales more favorably than G-POD. This allows us to include a larger number of basis functions in the POD basis without sacrificing computational efficiency. Hyper-reduction techniques, including energy-conserving ones, can also be employed to further sparsen the ROM operators [hyperreduction_1, hyperreduction_2, KLEIN2024112697].
@@ 4.3 Energy conservation of the ROM
To ensure stability of the ROMs, we aim to mimic the energy conservation property of the FOM, see (8). To investigate if the constructed ROMs satisfy this property we define the ROM energy $E_{r}^{\text{ROM}}$ as
@@
employing the product rule of differentiation. From the final expression, we conclude that a ROM based on Galerkin projection inherits energy conservation and stability from the FOM. This is true for both orthogonal and non-orthogonal projections. This means that not only the G-POD ROM inherits this property, as shown in [podbenjamin], but also the newly introduced space-local approaches L-POD and LO-POD.
@@ 5.1 Test case setup
To construct a ROM, we first require data from the FOM. As stated earlier, we use the finite difference discretization of the advection equation, detailed in Section 2.2, for this purpose. The system is simulated on a periodic domain $\Omega=[0,2\pi)$ discretized with $N=1000$ grid points for $t=[0,5]$ and constant $c=1$. A time step size of $\Delta t=0.01$ is used for the time integration, using a RK4 scheme. This is the largest time step size that still yielded stable simulations. Snapshot data for the ROM construction is collected in the interval $t=[0,1]$. We refer to this as the training data. Data generated in the interval $t=(1,2]$ will be referred to as validation data. This data is used to evaluate the generalization of the POD basis. The remaining part of the simulation data, $t=(2,5]$, will be used to evaluate the extrapolation capabilities of the ROMs. As an initial condition, we use a Gaussian wave centered around $x=\frac{1}{4}\pi$, namely
@@
Figure 5: Reference simulation of a Gaussian wave being advected throughout the domain. White bars separate the training data from the validation data and the validation data from the extrapolation data, respectively.
For the ROMs we consider the three bases discussed in this work: the global POD basis G-POD, the space-local POD basis L-POD, and the space-local POD basis with overlapping subdomains LO-POD. For the space-local approaches, the domain is subdivided into $I$ subdomains. For the local basis $\{\chi_{ij}\}$, we use $J$ box functions contained within each subdomain, such that $IJ=N$. In this way, the local basis aligns with the FOM finite difference basis, see (10). The integrals in the LO-POD post-processing step, equation (28), are approximated using the midpoint rule for integration [midpoint_dragomir1998some]. For the ROM time integration, we use the same RK4 scheme as for the FOM.
@@
[…]
where $\mathbf{P}$ projects the solution on the POD basis, see (31). This difference will be referred to as the solution error. In (41), this error is decomposed as the sum of the ROM error (the error made by the ROM during the time integration) and the projection error (the error made by the reduced basis approximation). Our implementation, in Julia [Julia-2017], of the introduced methodologies and experiments is freely available on Github, see https://github.com/tobyvg/local_POD_overlap.jl.